OpenStudy (aaronandyson):

A body is projected up reaches a point A in its path at the end of the 4th second and reaches the ground after 5 seconds from the rest.the height of A above the ground is?

OpenStudy (pythagoras123):

Can it be assumed the ball is at rest at point A?

OpenStudy (partyshotz2409):

y = -1/2 * 10*t^2 + u*t given t = 5 , y = 0 y(5) = 0 y(5) = -125 + 5u substitute -125 + 5u = 0 5u = 125 u = 125/5 u = 25 y = -1/2 * 10*t^2 + 25*t plug in t = 4 y(4) = 20 at 4 seconds the object is 20 meters above ground

OpenStudy (drmoss):

Alrighty young man I gave you the Best Response of the question